A company just paid dividends of $500,000.  The company had retained earnings of $1.5 mm.  500,000 shares are currently outstanding.  If the benchmark P/E is 16, what is the projected price of the stock?
Answer #1

Total earnings =1500000+500000 =2000000

Earnings per share =2000000/500000 =4

Projected price =16*4 =64